OpenStudy (anonymous):

5.71 grams of an unknown substance containing carbon, hydrogen and oxygen is burned in excess oxygen to generate 11.78 g of CO2 and 5.32 g of water. Determine the empirical formula for the unknown compound. Which is it? -C3H6O - C3H2O2 -C3H3O -CHO - C2H3O - CH2O

OpenStudy (triciaal):

write a balanced equation substance + oxygen generates carbon dioxide and water molecular mass of CO2 = 44 moles = mass/molecular mass = 11.78/44 = 0.268 molecular mass of water = 18 moles = 5.32/18 = 0.295 molecular mass of oxygen =32 x + 32 = 44 + 18 x = 30